# Inflection Point Inflation in Supergravity

### v1

@@ Abstract
In this paper, we study the inflection point inflation generated by a polynomial superpotential and a canonical Kähler potential under the supergravity framework, where only one chiral superfield is needed. We find that the special form of the scalar potential limits the inflationary Hubble parameter to values $\lesssim 10^{10}\,\textrm{GeV}$ and the inflaton mass to $\lesssim 10^{11}\,\textrm{GeV}$. We obtain analytic results for small field cases and present numerical results for large field ones. We find the tensor-to-scalar ratio $r<10^{-8}$ is always suppressed in these models, while the running of spectral index $\alpha\approx\mathcal{O}(-10^{-3})$ may be testable in next-generation CMB experiments. We also discuss the possible effects of SUSY breaking Polonyi term presented in the superpotential where we find a general upper bound for the SUSY breaking scale for a given value of the Hubble parameter.
@@ 1 Introduction
Recent observations of the cosmic microwave background (CMB) by the Planck and BICEP/Keck experiments strongly favor an exponentially expanding period of the early universe Planck:2018jri ; BICEP:2021xfz , which is called inflation Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . The simplest implementation, slow roll inflation, needs a single scalar field “inflaton” $\phi$ rolls over a flat region in the potential. The flatness of the potential can be parameterized by the following slow roll parameters [^1] We set the reduced Planck Mass $M_{\textrm{pl}}=\sqrt{\frac{1}{8\pi G}}\approx 2.4\times 10^{18}\,\textrm{GeV}=1$ throughout this paper. Lyth:2009zz :
$\epsilon=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}\,,\quad\eta=\frac{V^{\prime\prime}}{V}\,,\quad\xi^{2}=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}\,.$ (1)
where a prime () denotes the derivative with respect to $\phi$. They are related with the CMB observations Planck:2018jri ; BICEP:2021xfz :
@@
Here $P_{\zeta}$ is the power spectrum of Gaussian curvature perturbations, $n_{s}$ is the spectral index, $r$ is the tensor-to-scalar ratio and $\alpha$ is is the running of the spectral index.
@@
Power law potentials $V\propto\phi^{n}$ offer one of the most elegant and simplest realizations of inflation. However, such potentials with $n>1$ have been ruled out by the CMB observations, due to the large tensor-to-scalar ratio predicted by these models. Given this fact, a broad program has been pursed to investigate alternative models for inflation. Among them are the so-called inflection point inflation models, where different $\phi^{n}$ terms collectively provide an inflection point in the potential and accommodate inflation around it. Inflection point models have attracted much attention in recent years  Enqvist:2010vd ; Mazumdar:2011ih ; Hotchkiss:2011am ; Chang:2013cba ; Gao:2015yha ; Okada:2016ssd ; Gao:2016xfv ; Choi:2016eif ; Dimopoulos:2017xox ; Okada:2017cvy ; Pallis:2023aom . They have been well studied both during and after inflation Drees:2021wgd ; Drees:2022aea ; Drees:2024hok ; Bernal:2024ykj ; Drees:2025iue .
Loop corrections can easily spoil the flatness of the inflaton potential, while supersymmetry (SUSY) provides a convenient way to protect it. For this reason, different inflation models have often been considered in a supergravity (SUGRA) framework Nanopoulos:1982bv ; Ovrut:1983my ; Holman:1984yj ; Adams:1992bn ; Copeland:1994vg ; Kumekawa:1994gx ; Izawa:1996dv ; Linde:1997sj ; Binetruy:2004hh ; Allahverdi:2006we ; Ben-Dayan:2008fhy ; Ellis:2013xoa ; Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2014rga ; Ketov:2014qha ; Carrasco:2015pla ; Romao:2017uwa . This paper will follow the same spirit to consider a supersymmetric inflection point inflaton model. We want to present a systematic study allowing us to scan over the possible parameter space. We will focus on the similarities and differences between the previously studied non-supersymmetric, renormalizable models and SUGRA models.
SUSY must be broken since no superpartner of a Standard Model particle has as yet been detected. Once SUSY breaking terms are included, they may modify the form of the inflaton potential, thereby breaking the slow-roll conditions. Thus, the SUSY breaking scale must be bounded from above for successful inflation. To investigate this effect, we will focus on the case where SUSY is broken by the Polonyi model Polonyi:1977pj . We find two limiting cases by comparing the relative strength of the inflation sector and the SUSY breaking sector. They lead to very different bounds on the SUSY breaking scale.
@@
To be more precise, we consider a SUSY-preserving inflaton field $\phi$, which accommodates a near inflection point in the scalar potential at $\phi=\phi_{0}$. We find that the Hubble parameter and the inflaton mass both increase with $\phi_{0}$ when $\phi_{0}<1$. They reach a maximum around $\phi_{0}\sim 1$, then start to decrease because of the exponential factor in the potential. Thus, in such a model, the Hubble parameter can not exceed $\mathcal{O}(10^{10})\textrm{GeV}$. The tensor-to-scalar ratio $r$ is much smaller than the current upper bound. The model also predicts a near constant running $\alpha\sim-0.003$, a unique feature that the next generation of observations might be able to test.
@@
2 Renormalizable Inflection Point Model and Beyond
@@ 2.1 Potential setup and CMB observables
We start from the most general renormalizable potential of a single real scalar inflaton field $\phi$:
.$\,.V(\phi)=b\phi^{2}+c\phi^{3}+d\phi^{4}\,,$ (3)
@@
There are three free parameters in the potential, $d,\phi_{0}$ and $\beta$. In eq.(5), $d$ determines the overall normalization of the potential, which can be matched to the power spectrum of curvature perturbation $P_{\zeta}$ once the other parameters are fixed. The other two parameters govern the shape the potential and hence determine the CMB observables, such as the number of e-folds when the CMB pivot scale left the horizon, $N_{\textrm{cmb}}$, and the spectral index $n_{s}$. In a “small field” set–up, i.e. for sub-Planckian field values, $\phi_{0}<1$, fixing $N_{\textrm{cmb}}=65$ and eq.(2) require Drees:2021wgd :
@@ 2.2 Possible Realization in SUGRA and Associated Problem
In SUGRA the scalar potential is generated by the Kähler potential $K$ and the holomorphic Superpotential $W$ [^3] For a detailed introduction, please consult Freedman:2012zz .. Both of them are supposed to be functions of the complex field $\Phi$, which is the scalar component of a chiral superfield. The Kähler potential also determines the kinetic term of the scalar field:
@@
A canonically normalized field thus requires $K=-\Phi\Phi^{*}$ and we will use this Kähler potential throughout the paper. We can further define the Kähler covariant derivative as:
$D_{\Phi}W=\frac{\partial W}{\partial\Phi}+\frac{\partial K}{\partial\Phi}W\,.$ (9)
@@
where we neglect the constant and linear terms as in the renormalizable case. The ansatz (11) ensures that $V(0)=W(0)=0$, i.e. the potential has a supersymmetric stationary point at $\Phi=0$. The three coefficients $B,C,D$ are chosen to be real. The complex $\Phi$ can in general be written as $\Phi=(\phi+i\chi)/\sqrt{2}$, where $\phi$ and $\chi$ are real fields. We want to identify $\phi$ with the inflaton field. By choosing the coefficients in eq.(11) to be real we make sure that the potential $V(\phi,\chi)$ depends only on even powers of $\chi$. Moreover, we make sure that $\partial^{2}V(\phi,\chi)/\partial\chi^{2}>0$ for $\chi=0$ and $\phi\in[0,\phi_{0}]$. We therefore can assume that $\chi=0$ throughout, so that we can ignore this imaginary component when computing the inflationary dynamics. It is not hard to write down the scalar potential $V(\phi)\equiv V(\phi,\chi=0)$ in this setup:
$\begin{split}V(\phi)={\rm e}^{\frac{\phi^{2}}{2}}\Biggl{[}2B^{2}\phi^{2}+3\sqrt{2}BC\phi^{3}+\frac{1}{4}\phi^{4}\left(B^{2}+16BD+9C^{2}\right)+\frac{1}{2}\phi^{5}\left(\sqrt{2}BC+6\sqrt{2}CD\right)\\ \indent+\frac{1}{8}\phi^{6}\left(B^{2}+6BD+3C^{2}+16D^{2}\right)+\frac{1}{8}\phi^{7}\left(\sqrt{2}BC+4\sqrt{2}CD\right)\\ \indent+\frac{1}{16}\phi^{8}\left(2BD+C^{2}+5D^{2}\right)+\frac{CD\phi^{9}}{8\sqrt{2}}+\frac{D^{2}\phi^{10}}{32}\Biggr{]}\,,\end{split}$ (12)
In order to try to match to the renormalizable case, we expand $V$ up to the fourth order:
@@
Matching eq.(13) to eq.(3) and using eqs.(6) and (7) implies:
@@
Thus, when we evaluate the value of the potential and its first and second derivatives at $\phi\sim\phi_{0}$, the first five terms (from $\phi^{2}$ to $\phi^{6}$) would contribute with comparable magnitude. Therefore the terms $\propto\phi^{5}$ and $\phi^{6}$ can easily spoil the flatness of the potential, i.e. the expansion only to order $\phi^{4}$ is not self–consistent. For this reason, it is necessary to consider an inflection point model in the full potential eq.(12) (or at least up to $\phi^{6}$) rather than trying to directly match the renormalizable, non-supersymmetric potential.
@@ 3.1 Analytic Analysis of the Model
$\begin{split}B=&D\frac{\phi_{0}^{2}\left(1152+480\phi_{0}^{2}+72\phi_{0}^{4}+12\phi_{0}^{6}+\phi_{0}^{8}\right)}{2\left(192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}\right)}\\ C=&-D\frac{\sqrt{2}\phi_{0}\left(384+192\phi_{0}^{2}+24\phi_{0}^{4}+8\phi_{0}^{6}+\phi_{0}^{8}\right)}{192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}}\end{split}$ (17)
@@
The potential (18) has a minimum at $\phi=0$ and is positive semi-definite, i.e. $V(\phi)\geq 0\ \forall\phi$. As in the non-supersymmetric case, in the small field scenario $\phi_{0}\ll 1$ we can get semi-analytic results for inflationary observables by expanding the slow-roll parameter around the inflection point via the ansatz $\phi=\phi_{0}(1-\delta\phi)$, and adding a deviation from the strict inflection point condition via $B\to B+D\times\delta B$:
@@
We are now ready to discuss which $\delta\phi$ and $\delta B$ reproduce the CMB observations. In our model the duration of inflation is controlled by $\eta$ since $\epsilon\ll\eta$. The beginning (when the pivot scale $k_{*}=0.05\,\textrm{Mpc}^{-1}$ crossed out of the horizon) and end of observable inflation are given by:
@@
[…]
In the last step we have switched from the absolute deviation $\delta B$ to the relative one $\beta=D\times\delta B/B=\delta B/(3\phi_{0}^{2})$ in order to factor out $\phi_{0}$ in the denominator; except for the first line, eq.(23) also only holds for $\phi_{0}\ll 1$. Recall that the arc-tangent functions can be at most $\pi/2$. Numerically, requiring $N_{\textrm{cmb}}=50$ yields $\beta=2.7\times 10^{-5}\phi_{0}^{4}$ and $\delta B=8.2\times 10^{-5}\phi_{0}^{6}$. From the second eq.(2) and remembering $|\epsilon|\ll|\eta|$ we see that $|\eta|\geq 0.015$ when and after CMB scales crossed out of the horizon. Hence we can neglect the second term in eq.(22) for $\phi_{0}\ll 1$, and determine the inflation period by:
@@
$N_{\textrm{cmb}}=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left(\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{end}}}{12\sqrt{\beta}}}\right)-\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{cmb}}}{12\sqrt{\beta}}}\right)\right)\,,$ (25)
@@
The tensor-to-scalar ratio $r=16\epsilon$ is always too small to reach the sensitivity of any currently conceivable observation. However, the S4CMB experiment, together with small-scale structure information (e.g. on the Lyman-$\alpha$ forest) could achieve $10^{-3}$ sensitivity for $\alpha$, which would test our model Munoz:2016owz .
@@
By comparing with the predictions of the renormalizable model given in eqs.(7), we see that of $H_{\textrm{inf}}$ and $m_{\phi}$ scale with $\phi_{0}$ in the same way in both cases, with roughly a factor $2$ difference in coefficients. Even though the SUGRA potential is more complicated and scales as $\phi^{6}$ in the simplest limit, they thus make very similar predictions for the inflationary observables. In particular, the overall coefficient $D$ of the sixth order SUGRA potential is independent of $\phi_{0}$, while in the non-SUSY case the coefficient of the quartic potential $d\propto\phi_{0}^{2}$; hence in both cases $V(\phi\simeq\phi_{0})\propto\phi_{0}^{6}$.

### v2

@@ Abstract
In this paper, we study the inflection point inflation generated by a polynomial superpotential and a canonical Kähler potential under the supergravity framework, where only one chiral superfield is needed. We find that the special form of the scalar potential limits the inflationary Hubble parameter to values $\lesssim 10^{10}\,\textrm{GeV}$ and the inflaton mass to $\lesssim 10^{11}\,\textrm{GeV}$. We obtain analytic results for small field cases and present numerical results for large field ones. We find the tensor-to-scalar ratio $r<10^{-8}$ is always suppressed in these models, while the running of spectral index $\alpha\approx\mathcal{O}(-10^{-3})$ may be testable in next-generation CMB experiments. We also discuss the possible effects of a SUSY breaking Polonyi term presented in the superpotential where we find a general upper bound for the SUSY breaking scale for a given value of the Hubble parameter.
@@ 1 Introduction
Recent observations of the cosmic microwave background (CMB) by the Planck and BICEP/Keck experiments strongly favor an exponentially expanding period of the early universe Planck:2018jri ; BICEP:2021xfz , which is called inflation Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . The simplest implementation, slow roll inflation, needs a single scalar “inflaton” field $\phi$ rolling over a flat region in the potential. The flatness of the potential can be parameterized by the following slow roll parameters [^2] We set the reduced Planck Mass $M_{\textrm{pl}}=\sqrt{\frac{1}{8\pi G}}\approx 2.4\times 10^{18}\,\textrm{GeV}=1$ throughout this paper. Lyth:2009imm :
$\epsilon=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}\,,\quad\eta=\frac{V^{\prime\prime}}{V}\,,\quad\xi^{2}=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}\,$, (1)
where a prime () denotes the derivative with respect to $\phi$. They are related to the CMB observations Planck:2018jri ; BICEP:2021xfz : [^3] We will comment on the very recent ACT determination ACT:2025fju ; ACT:2025tim of $n_{s}$ in Sec. 3.1, following eq.(31).
@@
Here $P_{\zeta}$ is the power spectrum of Gaussian curvature perturbations, $n_{s}$ is the spectral index, $r$ is the tensor-to-scalar ratio and $\alpha$ is the running of the spectral index.
@@
Power law potentials $V\propto\phi^{n}$ offer one of the most elegant and simplest realizations of inflation. However, such potentials with $n>1$ have been ruled out by the CMB observations, due to the large tensor-to-scalar ratio predicted by these models. Given this fact, a broad program has been pursued to investigate alternative models for inflation. Among them are the so-called inflection point inflation models, where different $\phi^{n}$ terms collectively provide an inflection point in the potential and accommodate inflation around it. It has been realized using flat directions of the Minimally Supersymmetric Standard Model Allahverdi:2006iq ; Allahverdi:2006cx ; Allahverdi:2006we ; Enqvist:2010vd , in the context of Higgs inflation Okada:2016ssd ; Okada:2017cvy and warped $D$-brane inflation models Baumann:2007np ; Baumann:2007ah . As a generic model, it has been well studied both during and after inflation Hotchkiss:2011am ; Chang:2013cba ; Martin:2013tda ; Martin:2013nzq ; Choi:2016eif ; Dimopoulos:2017xox ; Musoke:2017frr ; Drees:2021wgd ; Drees:2022aea ; Drees:2024hok ; Bernal:2024ykj ; Drees:2025iue .
Supergravity emerges as the low energy limit of superstring theory, which is probably our best bet for a perturbative understanding of quantum gravity. Moreover, supersymmetry protects the (super)potential from loop corrections, which otherwise can easily spoil the required flatness of the inflaton potential. [^4] In large field models, quantum corrections to the Kähler metric can still be dangerous. In this paper we therefore consider inflation models in the supergravity (SUGRA) framework Nanopoulos:1982bv ; Adams:1992bn ; Copeland:1994vg ; Kumekawa:1994gx ; Izawa:1996dv ; Panagiotakopoulos:1997qd ; Linde:1997sj ; Kawasaki:2000yn ; Binetruy:2004hh ; Ellis:2013xoa ; Ellis:2013nxa ; Kallosh:2013yoa ; Carrasco:2015pla . Supergravity models with inflection point inflation have previously been constructed with canonical Kähler potential in Holman:1984yj ; Ross:1995dq ; Mazumdar:2011ih , with logarithmic Kähler potential in Linde:2007jn ; Conlon:2008cj ; Maharana:2015saa ; Gao:2015yha ; Pallis:2023aom , through a stabilizer field in Gao:2016xfv and with extended supergravity in Aldabergenov:2020bpt .
In this paper, we focus on a set–up with a single chiral superfield. A no-go theorem shows that a single chiral superfield with logarithmic Kähler potential can not support inflation Badziak:2008yg ; Ben-Dayan:2008fhy ; Covi:2008cn , unless a modification of Kähler potential is made Ketov:2014qha . Hence, we choose to work with canonical Kähler potential. For simplicity, we consider a polynomial superpotential with three terms. This mirrors non–supersymmetric inflection point inflation with renormalizable potential Drees:2021wgd , and allows us to control the position of the inflection point explicitly. Flexibility and control are among the novel features of our model.
There is a price we have to pay. Like many other inflection point models, parameters in this model are fine-tuned to support inflation and fit the CMB observables. There is no symmetry argument indicating these parameters should take specific values. In this sense, our model belongs to the category where inflation happens accidentally in the parameter space Linde:2007jn . Some progress has been made to reduce the fine-tuning, as in  Enqvist:2010vd ; Hotchkiss:2011am ; Dimopoulos:2017xox . It is beyond the scope of this paper to address this problem.
SUSY must be broken since no superpartner of a Standard Model particle has as yet been detected. Once SUSY breaking terms are included, they may modify the form of the inflaton potential, thereby breaking the slow-roll conditions. These corrections will be under control if the SUSY breaking scale is not too large. We investigate this issue for the simple case where SUSY is broken by the Polonyi model Polonyi:1977pj , which can play an important role in cosmology. In particular, the late time decay of Polonyi particles may adversely impact Big Bang nucleosynthesis (BBN); this is known as the “Polonyi problem” Ellis:1986zt ; Coughlan:1983ci . In F-term inflation models, this leads to a stringent constraint on the Hubble parameter during inflation Ibe:2006am . For D-term inflation, it has been studied in Addazi:2017ulg ; Aldabergenov:2017bjt ; Aldabergenov:2017hvp ; Abe:2018rnu . Here we focus on the impact of the Polonyi superpotential on the inflation potential. It has been shown that no-scale Starobinsky-like inflation is only compatible with a Polonyi sector if the SUSY breaking scale is smaller than $10^{3}\,\textrm{TeV}$ Romao:2017uwa . In this paper, we analyze two limiting cases by comparing the relative strengths of the inflation sector and the SUSY breaking sector. They lead to very different bounds on the SUSY breaking scale, which can be quite large in our model. This is another new result presented in this paper, which defuses the Polonyi problem, since the lifetime of Polonyi particles scales with the inverse third power of their mass; they will thus decay before the onset of BBN if their mass exceeds some tens of TeV.
@@
In this paper we present a systematic study allowing us to scan over the possible parameter space based on the position of the inflection point. We will focus on the similarities and differences between the previously studied non-supersymmetric, renormalizable models and SUGRA models. To be more precise, we consider a SUSY-preserving inflaton field $\phi$, which accommodates a near inflection point in the scalar potential at $\phi=\phi_{0}$. We find that the Hubble parameter and the inflaton mass both increase with $\phi_{0}$ as long as $\phi_{0}<1$. They reach a maximum around $\phi_{0}\sim 1$, then start to decrease because of the exponential factor in the potential. Thus, in such a model, the Hubble parameter can not exceed $\mathcal{O}(10^{10})\ \textrm{GeV}$. The tensor-to-scalar ratio $r$ is much smaller than the current upper bound. The model also predicts a near constant running $\alpha\sim-0.003$, a unique feature that the next generation of observations might be able to test Kohri:2013mxa ; CMB-S4:2016ple ; Munoz:2016owz ; Modak:2021zgb ; Easther:2021rdg ; Bahr-Kalus:2022prj .
@@
2 Non-supersymmetric Inflection Point Model and Beyond
@@ 2.1 Potential setup and CMB observables
We start from the most general renormalizable potential of a single real scalar inflaton field $\phi$ as a non-supersymmetric example:
$V(\phi)=b\phi^{2}+c\phi^{3}+d\phi^{4}\,$, (3)
@@
There are three free parameters in the potential, $d,\phi_{0}$ and $\beta$. In eq.(5), $d$ determines the overall normalization of the potential, which can be matched to the power spectrum of curvature perturbation $P_{\zeta}$ once the other parameters are fixed. The other two parameters govern the shape the potential and hence determine the CMB observables, such as the number of e-folds of inflation after the CMB pivot scale left the horizon, $N_{\textrm{cmb}}$, and the spectral index $n_{s}$. In a “small field” set–up, i.e. for sub-Planckian field values, $\phi_{0}<1$, fixing $N_{\textrm{cmb}}=65$ and eq.(2) require Drees:2021wgd :
@@ 2.2 Possible Realization in SUGRA and Associated Problem
In SUGRA [^6] For a detailed introduction, please consult Freedman:2012zz . the scalar potential is generated by the real Kähler potential $K$ and the holomorphic Superpotential $W$. Both are supposed to be functions of the complex field $\Phi$, which is the scalar component of a chiral superfield. The Kähler potential also determines the kinetic term of the scalar field (in the convention of ref. Freedman:2012zz ):
@@
A canonically normalized field thus requires $K=\Phi\Phi^{*}$ and we will use this Kähler potential throughout the paper. We can further define the Kähler covariant derivative as:
$D_{\Phi}W=\frac{\partial W}{\partial\Phi}+\frac{\partial K}{\partial\Phi}W\,=\frac{\partial W}{\partial\Phi}+\Phi^{*}W\,.$ (9)
@@
where we neglect the constant and linear terms as in the renormalizable case. The ansatz (11) ensures that $V(0)=W(0)=0$, i.e. the potential has a supersymmetric stationary point at $\Phi=0$. The three coefficients $B,C,D$ are chosen to be real. The complex $\Phi$ can in general be written as $\Phi=(\phi+i\chi)/\sqrt{2}$, where $\phi$ and $\chi$ are real fields. We want to identify $\phi$ with the inflaton field. By choosing the coefficients in eq.(11) to be real we make sure that the potential $V(\phi,\chi)$ depends only on even powers of $\chi$. Moreover, we make sure that $\partial^{2}V(\phi,\chi)/\partial\chi^{2}>0$ for $\chi=0$ and $\phi\in[0,\phi_{0}]$. We therefore can assume that $\chi=0$ throughout, so that we can ignore this imaginary component when computing the inflationary dynamics. It is not hard to write down the scalar potential $V(\phi)\equiv V(\phi,\chi=0)$ in this setup. We first notice that the two terms in eq.(10) read:
[…]
The combination of them leads to the full potential:
[…]
We first try to match to the non-supersymmetric case by expanding $V$ up to the fourth order:
@@
Naively these are the leading terms if $\phi\ll 1$. Matching eq.(14) to eq.(3) and using eqs.(6) and (7) implies:
@@
Thus, when we evaluate the value of the potential and its first and second derivatives at $\phi\sim\phi_{0}$, the first five terms (from $\phi^{2}$ to $\phi^{6}$) would contribute with comparable magnitude. Therefore the terms $\propto\phi^{5}$ and $\phi^{6}$ can easily spoil the flatness of the potential, i.e. the expansion only to order $\phi^{4}$ is not self–consistent. For this reason, it is necessary to consider an inflection point model in the full potential eq.(13) (or at least keeping terms up to $\phi^{6}$ in the small field case) rather than trying to directly match the renormalizable, non-supersymmetric potential.
@@ 3.1 Analytic Analysis of the Model
$\begin{split}B=&D\frac{\phi_{0}^{2}\left(1152+480\phi_{0}^{2}+72\phi_{0}^{4}+12\phi_{0}^{6}+\phi_{0}^{8}\right)}{2\left(192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}\right)}\,;\\ C=&-D\frac{\sqrt{2}\phi_{0}\left(384+192\phi_{0}^{2}+24\phi_{0}^{4}+8\phi_{0}^{6}+\phi_{0}^{8}\right)}{192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}}\,.\end{split}$ (18)
The inflection point condition allows us to bypass the $\eta$ problem for the canonical Kähler potential Copeland:1994vg ; Stewart:1994ts , which assumes that a single term in the second derivatives $V^{\prime\prime}$, arising from the exponential factor, dominates over the other possible ones. For a scalar potential accommodating an inflection point, this assumption doesn’t hold, and a flat potential appears only when we go beyond it. As already noted in the Introduction, this requires tuning of a priori unrelated parameters, in particular of the ratio $B/C$, as described by eqs.(18).
@@
The potential (19) has a minimum at $\phi=0$, with $V(0)=0$. The algebraic equation $D_{\phi}W\pm\sqrt{3}W=0$ has no real root besides $\phi=0$ when using eq.(18), which can be verified by checking the discriminant of the equation with additional combination of coefficients Rees:1922 . Hence the full potential of eq.(13) is positive semi-definite, i.e. $V(\phi)\geq 0\ \forall\phi$. [This is immediately obvious for the simplified potential of eq.(19).] As in the non-supersymmetric case, in the small field scenario $\phi_{0}\ll 1$ we can get semi-analytic results for inflationary observables by expanding the slow-roll parameter around the inflection point via the ansatz $\phi=\phi_{0}(1-\delta\phi)$, and adding a deviation from the strict inflection point condition via $B\to B+D\times\delta B$:
@@
We are now ready to discuss which $\delta\phi$ and $\delta B$ reproduce the CMB observations. In our model the duration of inflation is controlled by $\eta$ since $\epsilon\ll|\eta|$. The beginning (when the pivot scale $k_{*}=0.05\,\textrm{Mpc}^{-1}$ crossed out of the horizon) and end of observable inflation are given by:
@@
$\begin{split}N_{\textrm{cmb}}&=\int_{\phi_{\textrm{cmb}}}^{\phi_{\textrm{end}}}\frac{1}{\sqrt{2\epsilon}}d\phi\\ \indent&=\int_{\phi_{\textrm{cmb}}}^{\phi_{\textrm{end}}}\frac{\phi_{0}^{3}}{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}d\phi\\ \indent&=-\int_{\delta\phi_{\textrm{cmb}}}^{\delta\phi_{\textrm{end}}}\frac{\phi_{0}^{4}}{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}d\delta\phi\\ \indent&=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left[\arctan\left({\frac{\delta\phi_{\textrm{end}}}{\sqrt{\beta}}}\right)-\arctan\left({\frac{\delta\phi_{\textrm{cmb}}}{\sqrt{\beta}}}\right)\right]\,.\end{split}$ (24)
In the last step we have switched from the absolute deviation $\delta B$ to the relative one $\beta=D\times\delta B/B=\delta B/(3\phi_{0}^{2})$ in order to factor out $\phi_{0}^{2}$ in the denominator; except for the first line, eq.(24) also only holds for $\phi_{0}\ll 1$. Recall that the arc-tangent functions can be at most $\pi/2$. Numerically, requiring $N_{\textrm{cmb}}=50$ yields $\beta=2.7\times 10^{-5}\phi_{0}^{4}$ and $\delta B=8.2\times 10^{-5}\phi_{0}^{6}$. From the second eq.(2) and remembering $|\epsilon|\ll|\eta|$ we see that $|\eta|\geq 0.015$ when and after CMB scales crossed out of the horizon. Hence we can neglect the second term in eq.(23) for $\phi_{0}\ll 1$, and determine the inflation period by:
@@
$N_{\textrm{cmb}}=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left[\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{end}}}{12\sqrt{\beta}}}\right)-\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{cmb}}}{12\sqrt{\beta}}}\right)\right]\,$, (26)
@@
The tensor-to-scalar ratio $r=16\epsilon$ is always too small to reach the sensitivity of any currently conceivable observation. However, the S4CMB experiment, together with small-scale structure information (e.g. on the Lyman-$\alpha$ forest) could achieve $10^{-3}$ sensitivity for $\alpha$, which would test our model Kohri:2013mxa ; CMB-S4:2016ple ; Munoz:2016owz ; Modak:2021zgb ; Easther:2021rdg ; Bahr-Kalus:2022prj .
Recently, the ACT experiment measured a slightly larger spectral index when combining their results with DESI and Planck data, $n_{s}=0.9743\pm 0.0034$ DESI:2024mwx ; ACT:2025fju ; ACT:2025tim . Using this central value gives us slightly different predictions. Taking $N_{\textrm{cmb}}=45$ as an example, we have:
$\begin{split}\delta B=6.88\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=1.05\times 10^{-8}\phi_{0}^{6}\,,\\ D^{-2}=3.85\times 10^{14}\,&,\quad H_{\textrm{inf}}=4.16\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=3.06\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0035\,.\end{split}$ (32)
while taking $N_{\textrm{cmb}}=65$, we have:
$\begin{split}\delta B=2.83\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=2.02\times 10^{-9}\phi_{0}^{6}\,,\\ D^{-2}=1.99\times 10^{15}\,&,\quad H_{\textrm{inf}}=1.83\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=1.34\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0015\,.\end{split}$ (33)
One can see a shift in the spectral index leaves a subtle change on inflationary observables. The larger value of the spectral index from ACT would prolong the duration of inflation in the field space. Hence for a fixed number of e-folds $N_{\textrm{cmb}}$, ACT results will predict larger $\epsilon_{\textrm{cmb}}$, larger Hubble scale value $H_{\textrm{inf}}$ and inflaton mass $m_{\phi}$. This effect can be clearly seen by comparing the results in eq.(31) and (33). However, the differences between the model parameters derived from PLANCK and ACT observations with fixed $N_{\rm CMB}$ are much smaller than the variations of these parameters when $N_{\rm CMB}$ is varied between $45$ and $65$. We will therefore continue to use the PLANCK derived parameter values and estimate their uncertainty by considering these two values of $N_{\rm CMB}$.
Before moving to the next part, we would like to briefly mention the possibility of Ultra-slow-roll (USR) in our model, as inflection point inflation is a well-known example to study the USR process. To have a USR period, we need the kinetic energy of the inflation field to be much larger than its slow-roll kinetic energy Dimopoulos:2017ged . In order to have a stable USR, the second order derivative of the potential has to be positive: $V^{\prime\prime}>0$ Pattison:2018bct . The first condition is up to the initial conditions of the inflaton field dynamics. In our calculation, we have assumed the evolution starts from the slow-roll phase. This assumption can be relaxed to incorporate USR. The second condition is stricter. Throughout the observable inflation, our potential features a negative slow roll parameter $\eta$, i.e., $V^{\prime\prime}<0$. This excludes a stable USR phase.
@@
By comparing with the predictions of the renormalizable model given in eqs.(7), we see that $H_{\textrm{inf}}$ and $m_{\phi}$ scale with $\phi_{0}$ in the same way in both cases, with roughly a factor $2$ difference in coefficients. Even though the SUGRA potential is more complicated and scales as $\phi^{6}$ in the simplest limit, they thus make very similar predictions for the inflationary observables. In particular, the overall coefficient $D$ of the sixth order SUGRA potential is independent of $\phi_{0}$, while in the non-SUSY case the coefficient of the quartic potential $d\propto\phi_{0}^{2}$; hence in both cases $V(\phi\simeq\phi_{0})\propto\phi_{0}^{6}$.